at a baseball game sam bought 2 hamburgers and 1 order of fries for a toal of $3.75. Erica bought 1 hamburger and 2 orders of fries for a total of $3.00. what is the cost of one hamburger and the cost of one order of french fries?
Answer by stanbon(29625)

at a baseball game sam bought 2 hamburgers and 1 order of fries for a total of $3.75. Erica bought 1 hamburger and 2 orders of fries for a total of $3.00. what is the cost of one hamburger and the cost of one order of french fries?
--------------------
2h + f = 375 cents
h + 2f = 300 cents
----------------
Multiply the 2nd equation by 2 to get:
2h+4f = 600 cents
-----------------
Subtract the 1st equation from the last to get:
3f = 225
f = 75 cents (cost of an order of fries)
----------------
Substitute into h+2f = 300 to solve for "h"
h+2*75 = 300
h + 150 = 300
h = $1.50 (cost of one hamburger)
====================================
Cheers,
Stan H.